OpenStudy (anonymous):

Adding and Subtracting Rational Expressions? 1. (3/x) +(13/x-11) 2. (5/4x^2y)-y/14xz Please explain how/show work Thank you :)

OpenStudy (luigi0210):

Well we need to get common denominators first, so for the first question: \[\LARGE \frac{3}{x}+\frac{13}{x-11}\] We need common denominator of x(x-11), so we multiply: \[\LARGE (\frac{3}{x}*\frac{x-11}{x-11})+(\frac{13}{x-11}*\frac{x}{x})\] Now just combine and simplify, it'd be best to leave the denominators as they are and distribute the numerator: \[\LARGE \frac{3x-33}{x(x-11)}+\frac{13x}{x(x-11)}\] We get: \[\LARGE \frac{3x-33+13x}{x(x-11)}\] Simplified: \[\LARGE \frac{16x-33}{x(x-11)}\] Hope that made sense :)